OpenStudy (anonymous):

Proof: TanXCotX/Sec^2X

OpenStudy (mertsj):

\[\frac{\sin x}{\cos x}\times \frac{\cos x}{\sin x}\div \sec ^2x=1\times\frac{1}{\sec ^2x}=\cos ^2x\]